what is the hybridisation state  of i3-.if it is sp, then how can we predict that it is sp.

Grade:12

2 Answers

Saurabh Kumar
8 years ago
Consider the central I atom in [I-I-I]^-. It has seven valence electrons (group 17) two of which are used to for the I:I σ bonds; one electron is added for the charge:
# lp = 1/2[7 -2 +1] = 3 lp This is an AX2E3 system and is based on trigonal bipyramidal coordination (sp^3d hybridization). Lone pairs are found to adopt the equatorial sites (less steric repulsion) and so the terminal I atoms adopt the two axial sites. The [I3]^- ion is predicted and found to be linear.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Consider the central I atom in [I-I-I]^-.
It has seven valence electrons (group 17) two of which are used to for the I:I σ bonds; one electron is added for the charge:
# lp = 1/2[7 -2 +1] = 3 lp
This is an AX2E3 system and is based on trigonal bipyramidal coordination (sp^3d hybridization).
Lone pairs are found to adopt the equatorial sites (less steric repulsion) and so the terminal I atoms adopt the two axial sites.
The [I3]^- ion is predicted and found to be linear.

Thanks and Regards